OpenStudy (anonymous):

when a polynomial f(x)= x^3+ax^2-20x+24 is divided by x-1, the remainder is 7. Find the value of a. With the value of a, find the three linear factor s of f(x).

OpenStudy (anonymous):

i found the value of a which is 2 , but i cant get the the answer correctly for the following question.

hero (hero):

Interesting. How'd you figure out that the a = 2? Here's what I did: If x^3 + ax^2 - 20x + 24 is divided by x - 1, then it has remainder 7. Therefore x^3 + ax^2 -20x + 17 should divide evenly with x - 1 So we need to split ax^2 so that x - 1 divides evenly: x^3 + ax^2 - 20x + 17 = x^3 - x^2 + 3x^2 - 3x -17x + 17 = x^2(x - 1) + 3x(x - 1) - 17(x - 1) = (x - 1)(x^2 + 3x - 17) So obviously -x^2 + 3x^2 = 2x^2 which means a = 2

hero (hero):

Now that we know a = 2, we replace a with 2 in the original polynomial: f(x)= x^3+2x^2-20x+24 Then factor it. By factor theorem, f(2) = 0, which means x - 2 is a factor. Therefore: x^3 + 2x^2 - 20x + 24 = x^3 - 2x^2 + 4x^2 - 8x -12x + 24 = x^2(x - 2) + 4x(x - 2) - 12(x - 2) = (x - 2)(x^2 + 4x - 12) Finish factoring x^2 + 4x - 12 to find the remaining factors.

OpenStudy (anonymous):

the way i found a=2 is since x-1 is a factor of f(x) so f(1)=(1)^3+a(1)-20(1)+24 7=1+a-20+24 a=2

hero (hero):

Yep, that's the easiest way to find it.

OpenStudy (anonymous):

oh okay ,thanks for your help , but actually i am in trouble with this type of question , might have more practices on it.